OpenStudy (anonymous):

Solve for x: one over four (4x + 15) = 24 49 two hundred seven over four eighty one over four 9

OpenStudy (anonymous):

i have no idea what it is can you help me?

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (avengedslipknot):

yeah this is a bit confusing for me :/

OpenStudy (anonymous):

ikr i need @Michele_Laino

OpenStudy (anonymous):

please help @xMissAlyCatx

OpenStudy (xmissalycatx):

It is 81 over 4! :)

OpenStudy (anonymous):

Yes that is Corret 81 over 4.

OpenStudy (michele_laino):

here, we have to multiply both sides by \(4\), so I get this: \[4 \cdot \left( {\frac{1}{4}} \right) \cdot \left( {4x + 15} \right) = 4 \cdot 24\]